OpenStudy (anonymous):

solve by elimination or substitution: 1st problem) x^2-y^2=15 and x+y=1 2nd problem) 2y^2 +3x=33 and x+4y+7=0

OpenStudy (anonymous):

Well if you notice that x^2 -y^2 is (x+y)(x-y) and x+y = 1 than: 1*(x-y) = 15 x+y = 1 This should be easy to solve

OpenStudy (anonymous):

is 8 and -7 the answers?

OpenStudy (anonymous):

Yes thats right! :)